Prove that 2/n4-3 if and only off 4/n2+3


1
Expert's answer
2021-03-31T14:26:59-0400

Direct Proof:

Assume that "2|n^4-3." This means that there is some integer "y" such that "n^4-3=2y." Then "n^4=2y+3=2(y+1)+1."

Since "y+1" is an integer, we see that "n^4" is odd. 

Let "x\in \Z." Then "x^2" is even if and only if "x" is even.

Taking the contrapositive of both directions of the biconditional gives: "x^2" is odd if and only if "x" is odd.

Let "x=n^2." Then "n^4" is odd if and only if "n^2" is odd. Then "n^2"is odd if and only if "n" is odd.

Therefore there is an integer "k" such that "n=2k+1."

We write


"n^2+3=(2k+1)^2+3=4k^2+4k+4"

"=4(k^2+k+1)"

Since "k^2+k+1" is an integer, we see that "4|(n^2+3)."


Contrapositive:

Assume that "4\not| (n^2+3)." This means that for "k\in \Z," either

(i)"n^2+3=4k+1," (ii) "n^2+3=4k+2," or (iii) "n^2+3=4k+3."

If "n" is even then there is an integer "m" such that


"n^2+3=(2m)^2+3=4m^2+3"

which implies that we must be in case (iii).

If "n" is odd then there is an integer "m" such that


"n^2+3=(2m+1)^2+3=4m^2+4m+4"


"=4(m^2+m+1)"

In this case we see that "4|(n^2+3)."This is contrary to our assumption that "4\not| (n^2+3)" so under our assumption this case can never occur.

Thus we only have to consider case (iii) "n^2+3=4m+3."

Rearranging we can write "n^2=4k." Hence


"n^4-3=(4m)^2-3=16m^2-3"

"=2(8m^2-1)+1."

This means that the remainder is "1" when "n^4-3" is divided by "2."Therefore,

"2\not|(n^4-3)."


Therefore for "n\in \Z" "2|(n^4-3)" if and only if "4|(n^2+3)."